# The Gorini-Kossakowski-Sudarshan-Lindblad problem and the geometry of CP maps

### v4

@@ 1 Introduction
Although both Kraus decompositions and Lindblad parametrizations both involve can-be-written type claims, Lindblad parametrizers show that the latter is in a sense much simpler as there are uniform ways of extracting a suitable $(\Psi,G,H)$ parametrization. This is of importance when dealing with time-varying generators or perturbation problems.
The second part of the paper, starting in §7, is devoted to handling the infinite-dimensional (separable) case, bootstrapping from the finite-dimensional theory by way of a sequence of finite-dimensional approximations. After a brief orientation toward the separable setting in §7, the machinery for such approximations is developed in §§8 under the rubric filtration; and 9 derives a metric ($d$-metric) on a base space of a filtration with respect to which the unit ball is compact.
@@ 4.3 Further examples
4.3.1 When are Kraus sums equal?
Lemma 4.2.
Proof.
which is to say, $\tilde{U}^{\dagger}y=\tilde{x}$. ∎
Proposition 4.3.
$\sum_{i=1}^{n}{\bm{\pi}}(h_{i})=\sum_{i=1}^{n+m}{\bm{\pi}}(k_{j})$ if and only if $\tilde{h}=U{k}$ for some unitary matrix $U$, where $\tilde{h}$ is $h$ padded out with $m$ zero entries.
Proof.
Remark 4.1.
@@ 5.3 $L$-parametrization and the $L$-cone
$\begin{array}[]{ccccccl}\mathsf{HP}({\mathcal{H}})&\times&\mathrm{SA}(\mathcal{H})&\times&\mathrm{SA}(\mathcal{H})&\xrightarrow{\gamma}&\mathsf{HP}({\mathcal{H}})\\ (\Psi&,&G&,&H)&\mapsto&\Psi-{\bm{\theta}}(\operatorname{\mathbbm{1}}_{\mathcal{H}}+G+iH).\end{array}$ (5.7)
@@
Supposing $\Psi\in\mathsf{CP}(\mathcal{H})$, evidently $\gamma(\Psi,G,H)\in\mathsf{CP}(\mathcal{H})$. Therefore the curve $t\mapsto\gamma(t\Psi,tG,tH)$ is in $\mathsf{CP}(\mathcal{H})$ for $0\leq t$, so that $L(\Psi,G,H)\in\mathsf{cp}^{+}(\mathcal{H})$. Moreover, if $\Psi=0$, then the entire curve in $\mathsf{CP}(\mathcal{H})$, so $L(0,G,H)\in\mathsf{cp}(\mathcal{H})$.
@@ 7.1 Trace-class/bounded duality
$\mathcal{B}^{1}(\mathcal{H})\subset\mathcal{B}^{2}(\mathcal{H})\subset\mathfrak{B}(\mathcal{H}).$ (7.1)
@@
where $\{h_{i}\}$ is any complete orthonormal system in $\mathcal{H}$. For such an operator, the trace “norm” $\|A\|_{1}$ is equal to its trace, This definition extends to a general bounded operator $B$ by use of polar decomposition: if $B=MU$ with $M\geq 0$ and $U$ unitary, then $\|B\|_{1}{\kern 1.99997pt}:{\kern-5.0pt}={\kern 1.99997pt}\|M\|_{1}$. If $\|A\|_{1}<\infty$, then $A$ is said to be trace-class. In that case, its trace is unambiguously given by (7.2). $\mathcal{B}^{1}(\mathcal{H})$, the space of all trace-class operators, is a Banach space under this norm. $A$ is Hilbert-Schmidt if $A^{\dagger}A$ is trace class. On the Hilbert-Schmidt operators we have the by-now familiar inner product $\left\langle{A}\middle|{B}\right\rangle=\operatorname{\mathrm{Tr}}A^{\dagger}B$, and norm $\|A\|_{2}=\sqrt{\left\langle{A}\middle|{A}\right\rangle}$. $\mathfrak{B}(\mathcal{H})$ is the space of all bounded linear operators, with operator norm $\|A\|=\sup\{\|Ah\|\,:\,\|h\|=1\}$. It, too, is a Banach space. The Hilbert-Schmidt inner product $\left\langle{A}\middle|{B}\right\rangle$ does not extend to $\mathfrak{B}(\mathcal{H})$, but we can allow $A$ to be any bounded operator if we simultaneously restrict $B$ to be trace-class. This pairing identifies $\mathfrak{B}(\mathcal{H})$ with the Banach space dual of $\mathcal{B}^{1}(\mathcal{H})$ as $B\mapsto\left\langle{B}\middle|{\;\cdot\;}\right\rangle$. However, $\mathcal{B}^{1}(\mathcal{H})$ is not the dual space of $\mathfrak{B}(\mathcal{H})$, i.e., $\mathcal{B}^{1}(\mathcal{H})$ is not reflexive. That $\mathfrak{B}(\mathcal{H})$ is the dual space of $\mathcal{B}^{1}(\mathcal{H})$ means that the operator norm satisfies
@@
The superoperators which interest us are natively in $\mathfrak{B}(\mathcal{B}^{1}(\mathcal{H}))$. Such a $\Lambda\in\mathfrak{B}(\mathcal{B}^{1}(\mathcal{H}))$ has adjoint $\Lambda^{\dagger}\in\mathfrak{B}(\mathfrak{B}(\mathcal{H}))$ defined by
@@
Usually, one uses a pairing $\mathfrak{B}(\mathcal{H})\times\mathcal{B}^{1}(\mathcal{H})$ which is linear in both variables: $(B,\rho)\mapsto\langle B,\rho\rangle=\operatorname{\mathrm{Tr}}B\rho$. Then, one defines the dual of a superoperator as $\langle\Lambda^{*},\rho\rangle=\langle B,\Lambda\rho\rangle$. For a general Banach space in place of $\mathcal{B}^{1}(\mathcal{H})$, this would be the only choice. We are using instead a sesquilinear pairing in keeping with the rigged Hilbert space idea and for maximum continuity with the finite-dimensional theory.
@@ 7.2 Strong-operator topology
    $A_{n}\stackrel{{\scriptstyle SOT}}{{\longrightarrow}}A$ in $\mathfrak{B}(\mathcal{H})$ implies $A_{n},\cdot\;]_{\pm}\stackrel{{\scriptstyle SOT}}{{\longrightarrow}}[A,\cdot\;]_{\pm}$ in $\mathfrak{B}(\mathcal{B}^{1}(\mathcal{H}))$.
@@
(d). For any $v\in\mathcal{H}$, $A_{n}v\to Av$ implies that $A_{n}{\bm{\pi}}[v]\to A{\bm{\pi}}[v]$ in $\mathcal{B}^{1}(\mathcal{H})$ ($A_{n}{\bm{\pi}}[v]$ is just operator multiplication), and therefore $A_{n}\rho\to A\rho$ if $\rho\in\mathcal{B}^{1}(\mathcal{H})$ has finite rank. Since $\mathcal{B}^{\scriptstyle{fin}}{\mathcal{H}}$ is dense in $\mathcal{B}^{1}(\mathcal{H})$, the shadowing lemma implies $A_{n}\rho\to A\rho$ for any $\rho$. A similar argument shows $\rho A_{n}\to\rho A$. Now, just add or subtract. ∎
@@ 7.3 Weak-operator topology
Convergence in the weak-operator topology (WOT) is an even more permissive mode of convergence. A sequence $A_{n}\in\mathfrak{B}(\mathcal{H})$ converges in WOT if $\left\langle{w}\middle|{A_{n}v}\right\rangle$ converges for every $w,v\in\mathcal{H}$. By a double application of the Banach-Steinhaus theorem, the limits define a bounded operator. In this Hilbert space case, it is sufficient to check just diagonal matrix elements $\left\langle{v}\middle|{A_{n}v}\right\rangle$, as the others follow by polarization. The concept of WOT convergence also applies to superoperators $(\Gamma_{n})$ in $\mathfrak{B}(\mathcal{B}^{1}(\mathcal{H}))$ thus: $\left\langle{B}\middle|{\Gamma_{n}\rho}\right\rangle$ converges for every $B\in\mathfrak{B}(\mathcal{H})$, $\rho\in\mathcal{B}^{1}(\mathcal{H})$. Alternatively, we may discuss WOT in terms of subbasic neighborhoods, just as for SOT. $A\mapsto|\left\langle{w}\middle|{Av}\right\rangle|$ is a seminorm, corresponding to a “weak ball” $\mathrm{ball}_{W}(w,v)=\left\{A\in\mathfrak{B}(\mathcal{H})\,\colon\,|\left\langle{w}\middle|{Av}\right\rangle|<1\right\}$.
@@ 7.4 Meanings of old notations in the separable setting
Much of the notation used in the finite-dimensional setting transfers smoothly. The source of most of the clarifications needed arise from the fact that many of the occurences of $\mathfrak{B}(\mathcal{H})$” should be replaced by $\mathcal{B}^{1}(\mathcal{H})$”. In particular, the superoperators which interest us are natively in $\mathfrak{B}(\mathcal{B}^{1}(\mathcal{H}))$. Hence we have
$\mathsf{CP}(\mathcal{H},\mathcal{K})\subset\mathsf{Mon}(\mathcal{H},\mathcal{K})\subset\mathsf{HP}(\mathcal{H},\mathcal{K})\subset\mathfrak{B}(\mathcal{B}^{1}(\mathcal{H}),\mathcal{B}^{1}(\mathcal{K})).$ (7.8)
@@
just as before. $\mathrm{Pos}(\mathcal{H})$ is now ambiguous. We will use an appended “++”, i.e., $\mathcal{B}^{1}(\mathcal{H})^{+}$ and $\mathfrak{B}(\mathcal{H})^{+}$ are the positive elements in the indicated spaces.
@@ 8.1 Motivational sketch
Now, the $\mathcal{H}_{\alpha}$ are (finite-dimensional) Hilbert spaces in their own right, as well as subspaces of $\mathcal{H}$. We can take an operator $A$ on $\mathcal{H}_{\alpha}$ and regard it as a finite-rank operator on $\mathcal{H}$. $Ah$ is computed by applying $A$ as given to $P_{\alpha}h$ and then regarding the resulting vector as a member of $\mathcal{H}$. Conversely, given an operator $A$ on $\mathcal{H}$, we get one on $\mathcal{H}_{\alpha}$, as a projection $\hat{P}_{\alpha}A$, given as $(\hat{P}_{\alpha}A)h=P_{\alpha}(Ah)$. Thus, we get $\mathfrak{B}(\mathcal{H}_{1}),\mathfrak{B}(\mathcal{H}_{2}),\ldots$ as nested subspaces of $\mathfrak{B}(\mathcal{H})$, and $\mathcal{B}^{1}(\mathcal{H}_{1}),\mathcal{B}^{1}(\mathcal{H}_{2}),\ldots$ as nested subspaces of $\mathcal{B}^{1}(\mathcal{H})$. $\mathfrak{B}(\mathcal{H}_{\alpha})$ and $\mathcal{B}^{1}(\mathcal{H}_{\alpha})$ contain exactly the same linear operators; it is just their norms that are different.
We can copy the pattern at the next level, obtaining $\mathfrak{B}(\mathcal{B}^{1}(\mathcal{H}_{\alpha}))$ as a subspace of $\mathfrak{B}(\mathcal{B}^{1}(\mathcal{H}))$. This is really the level at which we want to apply these ideas. For, given $\Lambda$ in $\mathsf{HP}(\mathcal{B}^{1}(\mathcal{H}))$, we can project it into $\mathsf{HP}(\mathcal{B}^{1}(\mathcal{H}))$, apply the minimal Lindblad parametrizer to obtain a representation $Z_{\alpha}$ in $\mathsf{Lind\,{\mathcal{H}_{\alpha}}}$, and regard $Z_{\alpha}$ as an element of $\mathsf{Lind\,{\mathcal{H}}}$. Is it the case, perhaps, that the $Z_{\alpha}$ converge to a Lindblad representation of $\Lambda$? Not quite. It turns out that some subsequence does converge to a Lindblad parametrization of $\Lambda$, in an appropriate metric.

### v5

@@ 1 Introduction
Although Kraus decompositions and Lindblad parametrizations both involve can-be-written type claims, Lindblad parametrizers show that the latter is in a sense much simpler as there are uniform ways of extracting a suitable $(\Psi,G,H)$ parametrization. This is of importance when dealing with time-varying generators or perturbation problems.
The second part of the paper, starting in §7, is devoted to handling the infinite-dimensional (separable) case, bootstrapping from the finite-dimensional theory by way of a sequence of finite-dimensional approximations. After a brief orientation toward the separable setting in §7, the machinery for such approximations is developed in §8 under the rubric filtration; and §9 derives a metric ($d$-metric) on a base space of a filtration with respect to which the unit ball is compact.
@@ 4.2 Kraus decomposition is extremal decomposition
4.2.1 When are Kraus sums equal?
Lemma 4.2.
Proof.
which is to say, $\tilde{U}^{\dagger}y=\tilde{x}$. ∎
Proposition 4.3.
$\sum_{i=1}^{n}{\bm{\pi}}(h_{i})=\sum_{i=1}^{n+m}{\bm{\pi}}(k_{j})$, $m\geq 0$, if and only if $\tilde{h}=U{k}$ for some unitary matrix $U$, where $\tilde{h}$ is $h$ padded out with $m$ zero entries.
Proof.
Remark 4.1.
4.2.2 The set of extreme CP maps is closed
The titular claim will be used again when we prove an analog for the separable case, but not otherwise.
A CP $\Lambda$ has a Kraus decomposition $\Lambda=\sum_{i=1}^{m}{\bm{\theta}}(A_{i})$ with a minimal number, $m$, of summands. This is the Kraus rank of $\Lambda$. The correspondence of $\mathsf{CP}(\mathcal{H},\mathcal{K})$ with $\mathrm{Pos}(\mathfrak{B}(\mathcal{H},\mathcal{K}))$ gives easy access to this quantity.
First, we have a simple result for linear operators generally
Lemma 4.4.
Rank (denoted $\operatorname{rk}$) is a lower semicontinuous function on $\mathfrak{B}(\mathcal{H},\mathcal{K})$. That is, $\left\{A\in\mathfrak{B}(\mathcal{H},\mathcal{K})\,\colon\,\operatorname{rk}A\leq m\right\}$ is closed.
Proof.
$\operatorname{rk}A>m$ if and only if there are linearly independent $k_{1},\ldots,k_{m+1}$ such that $A^{\dagger}k_{i}\neq 0$ for every $i$. This condition is clearly stable under sufficiently small perturbation of $A$, i.e., $\left\{A\,\colon\,\operatorname{rk}A>m\right\}$ is open. ∎
Since the Kraus rank of $\Lambda$ is equal to the ordinary rank of $\mathfrak{J}\Lambda$, lower semicontinuity of Kraus rank follows immediately. In particular, the subset of $\mathsf{CP}(\mathcal{H},\mathcal{K})$ with Kraus rank at most one is closed. But this is exactly the extreme CP maps.
@@ 5.3 $L$-parametrization and the $L$-cone
$\begin{array}[]{ccccccl}\mathsf{HP}({\mathcal{H}})&\times&\mathrm{SA}(\mathcal{H})&\times&\mathrm{SA}(\mathcal{H})&\xrightarrow{\gamma}&\mathsf{HP}({\mathcal{H}})\\ (\Psi&,&G&,&H)&\mapsto&\Psi+{\bm{\theta}}(\operatorname{\mathbbm{1}}_{\mathcal{H}}-G-iH).\end{array}$ (5.7)
@@
Now, ${\bm{\theta}}(\operatorname{\mathbbm{1}}_{\mathcal{H}}-G-iH)$ is CP. Since a sum of CP maps is also such, $\gamma(\Psi,G,H)$ is CP as soon as $\Psi$ is so. Assuming that, the half-curve ($0\leq t$) $t\mapsto\gamma(t\Psi,tG,tH)$ is in $\mathsf{CP}(\mathcal{H})$, so that $L(\Psi,G,H)\in\mathsf{cp}^{+}(\mathcal{H})$. Moreover, if $\Psi=0$, then the entire curve in $\mathsf{CP}(\mathcal{H})$, so $L(0,G,H)\in\mathsf{cp}(\mathcal{H})$.
@@ 7.1 Trace-class/bounded duality
$\mathfrak{B}^{1}(\mathcal{H})\subset\mathfrak{B}^{2}(\mathcal{H})\subset\mathfrak{B}(\mathcal{H}).$ (7.1)
@@
where $\{h_{i}\}$ is any complete orthonormal system in $\mathcal{H}$. For such an operator, the trace “norm” $\|A\|_{1}$ is equal to its trace, This definition extends to a general bounded operator $B$ by use of polar decomposition: if $B=MU$ with $M\geq 0$ and $U$ unitary, then $\|B\|_{1}{\kern 1.99997pt}:{\kern-5.0pt}={\kern 1.99997pt}\|M\|_{1}$. If $\|A\|_{1}<\infty$, then $A$ is said to be trace-class. In that case, its trace is unambiguously given by (7.2). $\mathfrak{B}^{1}(\mathcal{H})$, the space of all trace-class operators, is a Banach space under this norm. $A$ is Hilbert-Schmidt if $A^{\dagger}A$ is trace class. On the Hilbert-Schmidt operators we have the by-now familiar inner product $\left\langle{A}\middle|{B}\right\rangle=\operatorname{\mathrm{Tr}}A^{\dagger}B$, and norm $\|A\|_{2}=\sqrt{\left\langle{A}\middle|{A}\right\rangle}$. $\mathfrak{B}(\mathcal{H})$ is the space of all bounded linear operators, with operator norm $\|A\|=\sup\{\|Ah\|\,:\,\|h\|=1\}$. It, too, is a Banach space. The Hilbert-Schmidt inner product $\left\langle{A}\middle|{B}\right\rangle$ does not extend to $\mathfrak{B}(\mathcal{H})$, but we can allow $A$ to be any bounded operator if we simultaneously restrict $B$ to be trace-class. This pairing identifies $\mathfrak{B}(\mathcal{H})$ with the Banach space dual of $\mathfrak{B}^{1}(\mathcal{H})$ as $B\mapsto\left\langle{B}\middle|{\;\cdot\;}\right\rangle$. However, $\mathfrak{B}^{1}(\mathcal{H})$ is not the dual space of $\mathfrak{B}(\mathcal{H})$, i.e., $\mathfrak{B}^{1}(\mathcal{H})$ is not reflexive. That $\mathfrak{B}(\mathcal{H})$ is the dual space of $\mathfrak{B}^{1}(\mathcal{H})$ means that the operator norm satisfies
@@
The superoperators which interest us are natively in $\mathfrak{B}(\mathfrak{B}^{1}(\mathcal{H}))$. Such a $\Lambda\in\mathfrak{B}(\mathfrak{B}^{1}(\mathcal{H}))$ has adjoint $\Lambda^{\dagger}\in\mathfrak{B}(\mathfrak{B}(\mathcal{H}))$ defined by
@@
Usually, one uses a pairing $\mathfrak{B}(\mathcal{H})\times\mathfrak{B}^{1}(\mathcal{H})$ which is linear in both variables: $(B,\rho)\mapsto\langle B,\rho\rangle=\operatorname{\mathrm{Tr}}B\rho$. Then, one defines the dual of a superoperator as $\langle\Lambda^{*},\rho\rangle=\langle B,\Lambda\rho\rangle$. For a general Banach space in place of $\mathfrak{B}^{1}(\mathcal{H})$, this would be the only choice. We are using instead a sesquilinear pairing in keeping with the rigged Hilbert space idea and for maximum continuity with the finite-dimensional theory.
@@ 7.2 Strong-operator topology
    $A_{n}\stackrel{{\scriptstyle SOT}}{{\longrightarrow}}A$ in $\mathfrak{B}(\mathcal{H})$ implies $A_{n},\cdot\;]_{\pm}\stackrel{{\scriptstyle SOT}}{{\longrightarrow}}[A,\cdot\;]_{\pm}$ in $\mathfrak{B}(\mathfrak{B}^{1}(\mathcal{H}))$.
@@
(d). For any $v\in\mathcal{H}$, $A_{n}v\to Av$ implies that $A_{n}{\bm{\pi}}[v]\to A{\bm{\pi}}[v]$ in $\mathfrak{B}^{1}(\mathcal{H})$ ($A_{n}{\bm{\pi}}[v]$ is just operator multiplication), and therefore $A_{n}\rho\to A\rho$ if $\rho\in\mathfrak{B}^{1}(\mathcal{H})$ has finite rank. Since $\mathcal{B}^{\scriptstyle{fin}}{\mathcal{H}}$ is dense in $\mathfrak{B}^{1}(\mathcal{H})$, the shadowing lemma implies $A_{n}\rho\to A\rho$ for any $\rho$. A similar argument shows $\rho A_{n}\to\rho A$. Now, just add or subtract. ∎
@@ 7.3 Weak-operator topology
Convergence in the weak-operator topology (WOT) is an even more permissive mode of convergence. A sequence $A_{n}\in\mathfrak{B}(\mathcal{H})$ converges in WOT if $\left\langle{w}\middle|{A_{n}v}\right\rangle$ converges for every $w,v\in\mathcal{H}$. By a double application of the Banach-Steinhaus theorem, the limits define a bounded operator. In this Hilbert space case, it is sufficient to check just diagonal matrix elements $\left\langle{v}\middle|{A_{n}v}\right\rangle$, as the others follow by polarization. The concept of WOT convergence also applies to superoperators $(\Gamma_{n})$ in $\mathfrak{B}(\mathfrak{B}^{1}(\mathcal{H}))$ thus: $\left\langle{B}\middle|{\Gamma_{n}\rho}\right\rangle$ converges for every $B\in\mathfrak{B}(\mathcal{H})$, $\rho\in\mathfrak{B}^{1}(\mathcal{H})$. Alternatively, we may discuss WOT in terms of subbasic neighborhoods, just as for SOT. $A\mapsto|\left\langle{w}\middle|{Av}\right\rangle|$ is a seminorm, corresponding to a “weak ball” $\mathrm{ball}_{W}(w,v)=\left\{A\in\mathfrak{B}(\mathcal{H})\,\colon\,|\left\langle{w}\middle|{Av}\right\rangle|<1\right\}$.
@@ 7.4 Meanings of old notations in the separable setting
Much of the notation used in the finite-dimensional setting transfers smoothly. The source of most of the clarifications needed arise from the fact that many of the occurences of $\mathfrak{B}(\mathcal{H})$” should be replaced by $\mathfrak{B}^{1}(\mathcal{H})$”. In particular, the superoperators which interest us are natively in $\mathfrak{B}(\mathfrak{B}^{1}(\mathcal{H}))$. Hence we have
$\mathsf{CP}(\mathcal{H},\mathcal{K})\subset\mathsf{Mon}(\mathcal{H},\mathcal{K})\subset\mathsf{HP}(\mathcal{H},\mathcal{K})\subset\mathfrak{B}(\mathfrak{B}^{1}(\mathcal{H}),\mathfrak{B}^{1}(\mathcal{K})).$ (7.8)
@@
just as before. $\mathrm{Pos}(\mathcal{H})$ is now ambiguous. We will use an appended “++”, i.e., $\mathfrak{B}^{1}(\mathcal{H})^{+}$ and $\mathfrak{B}(\mathcal{H})^{+}$ are the positive elements in the indicated spaces.
@@ 8.1 Motivational sketch
Now, the $\mathcal{H}_{\alpha}$ are (finite-dimensional) Hilbert spaces in their own right, as well as subspaces of $\mathcal{H}$. We can take an operator $A$ on $\mathcal{H}_{\alpha}$ and regard it as a finite-rank operator on $\mathcal{H}$. $Ah$ is computed by applying $A$ as given to $P_{\alpha}h$ and then regarding the resulting vector as a member of $\mathcal{H}$. Conversely, given an operator $A$ on $\mathcal{H}$, we get one on $\mathcal{H}_{\alpha}$, as a projection $\hat{P}_{\alpha}A$, given as $(\hat{P}_{\alpha}A)h=P_{\alpha}(Ah)$. Thus, we get $\mathfrak{B}(\mathcal{H}_{1}),\mathfrak{B}(\mathcal{H}_{2}),\ldots$ as nested subspaces of $\mathfrak{B}(\mathcal{H})$, and $\mathfrak{B}^{1}(\mathcal{H}_{1}),\mathfrak{B}^{1}(\mathcal{H}_{2}),\ldots$ as nested subspaces of $\mathfrak{B}^{1}(\mathcal{H})$. $\mathfrak{B}(\mathcal{H}_{\alpha})$ and $\mathfrak{B}^{1}(\mathcal{H}_{\alpha})$ contain exactly the same linear operators; it is just their norms that are different.
We can copy the pattern at the next level, obtaining $\mathfrak{B}(\mathfrak{B}^{1}(\mathcal{H}_{\alpha}))$ as a subspace of $\mathfrak{B}(\mathfrak{B}^{1}(\mathcal{H}))$. This is really the level at which we want to apply these ideas. For, given $\Lambda$ in $\mathsf{HP}(\mathfrak{B}^{1}(\mathcal{H}))$, we can project it into $\mathsf{HP}(\mathfrak{B}^{1}(\mathcal{H}))$, apply the minimal Lindblad parametrizer to obtain a representation $Z_{\alpha}$ in $\mathsf{Lind\,{\mathcal{H}_{\alpha}}}$, and regard $Z_{\alpha}$ as an element of $\mathsf{Lind\,{\mathcal{H}}}$. Is it the case, perhaps, that the $Z_{\alpha}$ converge to a Lindblad representation of $\Lambda$? Not quite. It turns out that some subsequence does converge to a Lindblad parametrization of $\Lambda$, in an appropriate metric.